# Galled Perfect Transfer Networks

### v2

@@ 23       Let $T$ be the tree whose set of non-trivial clades is $F$
    We know that a galled-completable tree, if one exists, must contain the clades $A\setminus B,A\cap B$, or $B\setminus A,B\cap B$. We do not know which one it is, so we try the former first. That is, first consider the partition $\{A_{1},A_{2}\}=\{A\setminus B,A\cap B\}$.
@@
    Do note that Lemma 4 only applies to maximal characters. Since we potentially check them on $A$ and $B$, it is crucial to have both $A$ and $B$ as maximal characters.
@@ 5 Galled PTNs on functional characters: a case study
Figure 12: (a) The species tree with 45 species with annotated leaves showing colored annotations by clade according to the presence of one of the characters: $\{K13628,K19115,K00561,K19310\}$. Note that between annotations of the same color there should exist a transfer edge that could join the parts of the tree where the characters are present. Leaf names correspond to the species id present in the KEGG orthology database. Notice that the redundancy-free network is not galled since the cycles resulting from joining the first appearances of characters $K13628$ and $K00561$ contain the root of the tree as a common node. However as shown in (b), by resolving one of the children of the root node, as the black dot, this network becomes galled. Tree annotations where visualized using iTOL tool  [itol].
@@
Going back to Figure 10, we note that there do exist characters that possess at most two first-appearances. We checked whether this subset of four characters could at least be galled-completable, but it turns out that this is not even the case. This is shown in Figure 12 on the left. As explained in the caption, if we built the redundancy-free network for these characters and applied Lemma 3, we would add a transfer edge linking the branches above the roots of the blue character, and above the roots of the green character. These are non-redundant edges that create two cycles that intersect at the root. But one could hypothesize that the unresolved root is responsible for this: Figure 12 on the right partially resolves the tree by expanding the root, creating a branch in a way that the cycle of the blue characters now avoids the root. One can now check that the resulting redundancy-free network is galled, and that this set of characters is galled-completable.
@@ 6 Conclusion
In this work, we expanded the foundations on perfect phylogenies to galled trees. While compatible character in trees enjoy an elegant mathematical structure, it appears that the difficulty of characterizing compatibility ramps up very quickly as we move beyond trees. Nonetheless, this work can serve as a stepping-stone towards the understanding of character evolution on more complex structures. A small step in this direction would be to allow bidirectional transfers in galled trees. This change is not as trivial as it seems, since several of our key results do not hold when such edges are present. It will also be interesting to integrate broader classes of networks, for example level-$k$ networks, and to solve the open problem of finding a PTN with a minimum number of transfer edges.

### v3

@@ 23       Let $T$ be the tree whose set of non-trivial clades is $F$
    We know that a galled-completable tree, if one exists, must contain the clades $A\setminus B,A\cap B$, or $B\setminus A,A\cap B$. We do not know which one it is, so we try the former first. That is, first consider the partition $\{A_{1},A_{2}\}=\{A\setminus B,A\cap B\}$.
@@
    Do note that Lemma 4 only applies to maximal characters. Since we potentially check the four conditions of the lemma on $A$ both and $B$, it is crucial to have both $A$ and $B$ as maximal characters.
@@ 5 Galled PTNs on functional characters: a case study
Figure 12: (a) The species tree with 45 species with annotated leaves showing colored annotations by clade according to the presence of one of the characters: $\{K13628,K19115,K00561,K19310\}$. Note that between annotations of the same color there should exist a transfer edge that could join the parts of the tree where the characters are present. Leaf names correspond to the species id present in the KEGG orthology database. Notice that the resulting redundancy-free network is not galled since the cycles resulting from joining the first appearances of characters $K13628$ and $K00561$ contain the root of the tree as a common node. However as shown in (b), by resolving one of the children of the root node, as the black dot, this network becomes galled. Tree annotations where visualized using iTOL tool  [itol].
@@
Going back to Figure 10, we note that there do exist characters that possess at most two first-appearances. We checked whether this subset of four characters could at least be galled-completable, but it turns out that this is not even the case. This is shown in Figure 12 on the left. As explained in the caption, if we built a redundancy-free network for these characters and applied Lemma 3, we would add a transfer edge linking the branches above the roots of the blue character, and above the roots of the green character. These are non-redundant edges that create two cycles that intersect at the root. But one could hypothesize that the unresolved root is responsible for this: Figure 12 on the right partially resolves the tree by expanding the root, creating a branch in a way that the cycle of the blue characters now avoids the root. One can now check that the resulting redundancy-free network is galled, and that this set of characters is galled-completable.
@@ 6 Conclusion
In this work, we expanded the foundations on perfect phylogenies to galled perfect transfer networks. While compatible characters in trees enjoy an elegant mathematical structure, it appears that the difficulty of characterizing compatibility ramps up very quickly as we move beyond trees. Nonetheless, this work can serve as a stepping-stone towards the understanding of character evolution on more complex structures. A small step in this direction would be to allow bidirectional transfers in galled trees. This change is not as trivial as it seems, since several of our key results do not hold when such edges are present. It will also be interesting to integrate broader classes of networks, for example level-$k$ networks, and to solve the open problem of finding a PTN with a minimum number of transfer edges. In that regards, it is possible that the algorithms from [vanIersel2010, kelk2012] for explaining softwired clusters could be adapted to PTN, as only differ by restricting which reticulation edges can be switched off or not.